### v4

@@ 7.3 New Benchmark on SoTA Methods
For Object. Persona includes 47 subjects from available resources [7, 4, 20]. Following the methodology of DreamBooth [4], we categorize the subjects into two classes: objects and live pets, based on whether the subject is a living entity. Specifically, 10 out of 47 subjects are pets and the remaining 37 are various objects. To evaluate the performance, we utilize the text prompts from DreamBooth [4]. This includes 20 recontextualization prompts and 5 property modification prompts for objects, along with 10 recontextualization, 10 accessorization, and 5 property modification prompts for pets, totally 25 prompts per category.
For Face. We also collected 15 subjects from Celeb-A [159] into our Persona dataset. We use 40 prompts for evaluation, including 10 accessory prompts, 10 style prompts, 10 action prompts, and 10 context prompts.
Settings. To assess existing representative PCS methods using our constructed test dataset, we compute CLIP-T to assess the text alignment. We calculate CLIP-I for subject fidelity in object generation. For subject fidelity in face generation, we detect faces in both the generated and target images using MTCNN [160] and calculate the pairwise identity similarity using FaceNet [161]. To uniformly evaluate and compare existing personalized generation methods, we select 22 representative PCS methods for evaluation. We generate 4 images for each test prompt and set the same random seed for all methods.
@@
Object Textual Inversion [7] TTF SD 1.5 0.199 0.749
Dreambooth [4] TTF SD 1.5 0.286 0.772
P+ [18] TTF SD 1.4 0.244 0.643
Custom Diffusion [20] TTF SD 1.4 0.307 0.722
NeTI [30] TTF SD 1.4 0.283 0.801
SVDiff [22] TTF SD 1.5 0.282 0.776
Perfusion [21] TTF SD 1.5 0.273 0.691
ELITE [47] PTA SD 1.4 0.292 0.765
BLIP-Diffusion [57] PTA SD 1.5 0.292 0.772
IP-Adapter [51] PTA SD 1.5 0.2722 0.825
SSR Encoder [55] PTA SD 1.5 0.288 0.792
MoMA [59] PTA SD 1.5 0.322 0.748
Diptych Prompting [101] PTA FLUX 1.0 dev 0.327 0.722
$\lambda$-eclipse [112] PTA Kandinsky 2.2 0.272 0.824
MS-Diffusion [56] PTA SDXL 0.298 0.777
Face CrossInitialization [33] TTF SD 2.1 0.261 0.469
Face2Diffusion [162] PTA SD 1.4 0.265 0.588
SSR Encoder [55] PTA SD 1.5 0.233 0.490
FastComposer [5] PTA SD 1.5 0.230 0.516
IP-Adapter [51] PTA SD 1.5 0.292 0.462
IP-Adapter [51] PTA SDXL 0.292 0.642
PhotoMaker [46] PTA SDXL 0.311 0.547
InstantID [6] PTA SDXL 0.278 0.707
@@ 8.1 Overfitting Problem
To investigate the rationale behind, Compositional Inversion [87] observes that the learned token embedding is located in an out-of-distribution area compared to the center distribution formed by pre-trained words. This is also found in another work [33] that the learnable token embeddings deviate significantly from the distribution of the initial embedding. In addition, there is evidence [87, 76, 71] suggesting that the unique modifier dominates in the cross-attention layers compared to the other context tokens, leading to the absence of other semantic appearances.
To address this issue, many solutions have been proposed. Most methods discussed in Section 4 contribute to the alleviation of the overfitting problem, such as the exclusion of redundant background, attention manipulation, regularization of the learnable parameters, and data augmentation. However, it has not been solved yet, especially in the cases where the SoI has a non-rigid appearance [87] or the context prompt has a similar semantic correlation with the irrelevant elements in the reference [19]. It is clear that addressing overfitting in PCS is not merely a technical challenge but a necessity for ensuring the practical deployment and scalability of these systems in varied and dynamic real-world environments. Therefore, an effective strategy and robust evaluation metrics are urgently needed to achieve broader adoption and greater satisfaction in practical applications.
@@ 8.2 Trade-off on Subject Fidelity and Text Alignment
The ultimate goal of personalized content synthesis is to create systems that not only render the SoI with high fidelity but also effectively respond to textual prompts. However, achieving excellence in both areas simultaneously presents a notable conflict. Specifically, high subject fidelity typically involves capturing and reproducing detailed and specific features of the SoI. This often requires the model to minimize the reconstruction loss to replicate delicate characteristics accurately. Conversely, text alignment necessitates that the system flexibly adapts the SoI according to varying textual descriptions. These descriptions may suggest changes in pose, expression, environment, or stylistic alterations that do not aim to reconstruct the exact visualization in the reference. As a result, it becomes challenging to achieve flexible adaptation in different contexts while simultaneously pushing the model to capture fine-grained details. To address this inherent conflict, Perfusion [21] proposes to regularize the attention projections by these two items. [163] decouples the conditional guidance into two separate processes, which allows for the distinct handling of subject fidelity and textual alignment. Despite these efforts, there still remains room for further exploration and refinement of this issue. Enhanced model architectures, innovative training methodologies, and more dynamic data handling strategies could potentially provide new pathways to better balance the demands of subject and text fidelity in PCS systems.
@@ 8.4 Multimodal Autoregressive Frameworks
Recent advancements in multimodal autoregressive models present novel solutions for PCS by unifying cross-modal understanding and generation. Models like Emu3 [164] demonstrate that autoregressive architectures can natively handle image-text-video sequences through discrete tokenization and joint transformer training. This paradigm enables seamless integration of user-provided multimodal references (e.g., text descriptions paired with SoI images) while maintaining contextual coherence across generation steps. Besides, this framework natively support for subject editing via multi-round chat, effectively addressing the overfitting limitations commonly observed in diffusion-based models.
@@ 8.5 Interactive Personalization Workflow
The evolution of interactive generation systems has unlocked new frontiers for PCS, particularly through the integration of multi-round interactive generation. This capability allows users to iteratively refine and accurately define the SoI, addressing the challenge of translating vague or complex requirements into precise content generation. For instance, conversational PCS systems like Gemini-2.0-flash [165] exemplify this progress, leveraging natural language dialogue to iteratively optimize both subject fidelity and prompt alignment. By enabling users to provide real-time feedback and adjust parameters through chat-like interactions, these systems bridge the gap between abstract intent and concrete outputs, aligning with PCS’s core objective of balancing faithful subject representation with flexible editability.
@@ 9 Conclusion
Textual Inversion [7] Object TTF LDM Token Embedding Fine-tuning
@@
DreamBooth [4] Object TTF Imagen Diffusion Model Fine-tuning; Regularization Dataset
DreamArtist [92] Object TTF LDM Negative Prompt Fine-tuning
DVAR [90] Object TTF SD 1.5 Randomness Erasing
HiPer [166] Object TTF SD Token Embedding Enhancement
P+ [18] Object TTF SD 1.4 Token Embedding Enhancement
@@
Unet-finetune [23] Object TTF SD Parameter-efficient Fine-tuning
Jia et al. [76] Object TTF Imagen Mask-assisted Generation; Regularization
COTI [83] Object TTF SD 2.0 Data Augmentation
Gradient-Free TI [94] Object TTF SD Evolutionary Strategy
PerSAM [44] Object TTF SD Mask-assisted Generation
DisenBooth [77] Object TTF SD 2.1 Regularization
NeTI [30] Object TTF SD 1.4 Token Embedding Enhancement
Prospect [31] Object TTF SD 1.4 Token Embedding Enhancement
@@
Break-a-Scene [38] Object TTF SD 2.1 Attention-based Operation; Attention-based Operation; Mask-assisted Generation
COMCAT [96] Object TTF SD 1.4 Attention-based Operation
ViCo [71] Object TTF SD Attention-based Operation; Regularization
OFT [167] Object TTF SD 1.5 Finetuning Strategy
LyCORIS [36] Object TTF SD 1.5 Attention-based Operation
He et al. [40] Object TTF SD, SDXL Data Augmentation
DIFFNAT [168] Object TTF SD Regularization
@@
MATTE [91] Object; Style; High-level Semantic TTF SD Regularization
LEGO [116] Object TTF SD Regularization
Catversion [95] Object TTF SD 1.5 Embedding Concatenation
@@
CLiC [45] Object TTF SD 1.4 Attention-based Operation; Mask-assisted Generation
@@
HiFi Tuner [32] Object TTF SD 1.4 Attention-based Operation; Mask-assisted Generation; Regularization
InstructBooth [93] Object TTF SD 1.5 Reinforcement Learning
DETEX [49] Object TTF SD 1.4 Mask-assisted Generation
DreamDistribution [169] Object TTF SD 2.1 Token Embedding Enhancement
DreamTuner [39] Object TTF SD Attention-based Operation
Xie et al. [34] Object TTF SD Regularization; Mask-assisted Generation
Infusion [170] Object TTF SD 1.5 Attention-based Operation
@@
Pair Customization [171] Object TTF SDXL Disentanglement Approach
DisenDreamer [172] Object TTF SD 2.1 Disentanglement Approach
DreamBooth+++++ + [173] Object TTF SD Attention-based Operation
CoRe [174] Object TTF SD 2.1 Regularization
DCI [175] Object TTF SD Token Embedding Optimization
Re-Imagen [97] Object PTA Imagen Retrieval-augmented Paradigm
Versatile Diffusion [176] Object PTA SD 1.4 Architecture Design
Tuning-Encoder [88] Object PTA SD Regularization
@@
ELITE [47] Object PTA SD 1.4 Attention-based Operation; Mask-assisted Generation
UMM-Diffusion [61] Object PTA SD 1.5 Reference Feature Injection
SuTI [52] Object PTA Imagen Data Augmentation
InstantBooth [98] Object PTA SD 1.4 Patch Feature Extraction
@@
BLIP-Diffusion [57] Object PTA SD 1.5 Data Augmentation; Mask-assisted Generation
@@
Domain-Agnostic [89] Object PTA SD Regularization; Attention-based Operation
IP-Adapter [51] Object PTA SD 1.5 Reference Feature Injection
Kosmos-G [60] Object; Face PTA SD 1.5 Multimodal Language Modeling
Kim et al. [177] Object PTA SD 1.4 Reference Feature Injection
CAFÉ [58] Object; Face PTA SD 2.0 Multimodal Language Modeling
SAG [178] Object; Face; Style PTA SD 1.5 Gradient-based Guidance
BootPIG [179] Object PTA SD Data Augmentation
CustomContrast [180] Object PTA SD 1.5 Contrastive Learning
MoMA [59] Object PTA SD 1.5 Multimodal LLM Adapter
@@
Diptych Prompting [101] Object; Style PTA FLUX Attention-based Operation; Mask-assisted Generation
@@
StyleDrop [24] Style; Multiple Subjects TTF Muse Data Augmentation
StyleBoost [84] Style TTF SD 1.5 Regularization Dataset
StyleAligned [99] Style TTF SDXL Regularization
GAL [19] Style; Object TTF SD 1.5 Data Augmentation
@@
StyleForge [181] Style TTF SD 1.5 Parameter-efficient Fine-tuning
FineStyle [182] Style TTF Muse Mask-assisted Generation
Style-Friendly [100] Style TTF FLUX-dev; SD 3.5 SNR Sampler
@@
StyleAdapter [82] Style PTA SD Data Augmentation; Reference Feature Fxtraction
ArtAdapter [183] Style PTA SD 1.5 Data Augmentation
ProFusion [106] Face TTF SD 2.0 Fusion Sampling
Celeb Basis [65] Face TTF SD Face Feature Representation
Banerjee et al. [184] Face TTF SD 1.4 Regularization
@@
Magicapture [37] Face; Multiple Subjects TTF SD 1.5 Attention-based Operation; Mask-assisted Generation; Identity Loss
Concept-centric [163] Face TTF SD 1.5 Modified Classifier-free Guidance
Cross Initialization [33] Face TTF SD 2.1 Regularization
@@
OMG [27] Face; Multiple Subjects TTF SDXL Attention-based Operation; Mask-assisted Generation; Identity Loss
@@
InstantFamily [185] Face; Multiple Subjects; Extra Conditions TTF SD 1.5 Mask-assisted Generation
PersonaMagic [186] Face TTF SD 1.4 Attention-based Operation
HyperDreamBooth [26] Face TTF; PTA SD 1.5 Pretraining and Fast Fine-tuning
Su et al. [187] Face PTA LDM; StyleGAN Identity Loss; Multi-task Learning
@@
FastComposer [5] Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
Face0 [107] Face PTA SD 1.4 Fusion Sampling
DreamIdentity [48] Face PTA SD 2.1 Attention-based Operation
@@
Face-Diffuser [79] Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
W-Plus-Adapter [53] Face PTA SD 1.5; StyleGAN Face Feature Representation
Portrait Diffusion [188] Face PTA SD 1.5 Mask-assisted Generation
RetriNet [74] Face PTA SD Mask-assisted Generation
FaceStudio [109] Face PTA SD Attention-based Operation
PVA [189] Face PTA LDM Mask-assisted Generation
DemoCaricature [121] Face; Extra Conditions PTA SD 1.5 Regularization
@@
PhotoMaker [46] Face PTA SDXL Data Collection; Reference Feature Extraction
Stellar [35] Face PTA SDXL Evaluation Prompts and Metrics
@@
PortraitBooth [50] Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
@@
InstantID [6] Face; Extra Conditions PTA SDXL Multi-feature Injection
ID-Aligner [190] Face PTA SD 1.5; SDXL Feedback Learning
@@
MoA [191] Face; Multiple Subjects PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
IDAdapter [192] Face PTA SD 2.1 Mixed Facial Features
Infinite-ID [193] Face PTA SDXL Reference Feature Injection
Face2Diffusion [162] Face PTA SD Multi-feature Injection
FreeCure [194] Face PTA SD 1.5; SDXL Mask-assisted Generation
Omni-ID [195] Face PTA FLUX Face representation enhancement
@@
Custom Diffusion [20] Multiple Subjects TTF SD 1.4 Parameter-efficient Fine-tuning; Constrained Optimization
Cones [111] Multiple Subjects TTF SD 1.4 Concept Neurons Activation
@@
SVDiff [22] Multiple Subjects TTF SD Data Augmentation; Attention-based Operation
Perfusion [21] Multiple Subjects TTF SD 1.5 Regularization
@@
Mix-of-Show [25] Multiple Subjects TTF SD 1.5 Attention-based Operation; Mask-assisted Generation
@@
Cones-2 [70] Multiple Subjects TTF SD 2.1 Attention-based Operation; Regularization; Mask-assisted Generation
PACGen [75] Multiple Subjects TTF SD 1.4 Mask-assisted Generation
@@
Compositional Inversion [87] Multiple Subjects TTF SD Attention-based Operation; Regularization
@@
EM-style [72] Multiple Subjects TTF SD Attention-based Operation; Mask-assisted Generation
MC2 [196] Multiple Subjects TTF SD 1.5 Attention-based Operation
MultiBooth [64] Multiple Subjects TTF SD 1.5 Mask-assisted Generation
Matsuda et al. [197] Multiple Subjects TTF LDM Mask-assisted Generation
@@
MagicTailor [198] Multiple Subjects TTF SD 2.1 Mask-assisted Generation; Attention-based Operation
AnyDoor [78] Multiple Subjects PTA SD 2.1 Mask-assisted Generation
@@
Subject-Diffusion [63] Multiple Subjects PTA SD 2.0 Attention-based Operation; Data Augmentation; Mask-assisted Generation
@@
CustomNet [113] Multiple Subjects; Extra Conditions PTA SD Multi-condition Integration
@@
MIGC [114] Multiple Subjects PTA SD Attention-based Operation; Mask-assisted Generation
Emu2 [199] Multiple Subjects PTA SDXL In-Context Learning
SSR-Encoder [55] Multiple Subjects PTA SD 1.5 Attention-based Operation

### v5

@@ 7.3 New Benchmark on SoTA Methods
For Object. Persona includes 47 subjects from available resources [1gal2022image, 2ruiz2023dreambooth, 6kumari2023multi]. Following the methodology of DreamBooth [2ruiz2023dreambooth], we categorize the subjects into two classes: objects and live pets, based on whether the subject is a living entity. Specifically, 10 out of 47 subjects are pets and the remaining 37 are various objects. To evaluate the performance, we utilize the text prompts from DreamBooth [2ruiz2023dreambooth]. This includes 20 recontextualization prompts and 5 property modification prompts for objects, along with 10 recontextualization, 10 accessorization, and 5 property modification prompts for pets, totally 25 prompts per category.
For Face. We also collected 15 subjects from Celeb-A [liu2018large] into our Persona dataset. We use 40 prompts for evaluation, including 10 accessory prompts, 10 style prompts, 10 action prompts, and 10 context prompts.
Settings. To assess existing representative PCS methods using our constructed test dataset, we compute CLIP-T to assess the text alignment. We calculate CLIP-I for subject fidelity in object generation. For subject fidelity in face generation, we detect faces in both the generated and target images using MTCNN [zhang2016mtcnn] and calculate the pairwise identity similarity using FaceNet [schroff2015facenet]. To uniformly evaluate and compare existing personalized generation methods, we select 22 representative PCS methods for evaluation. We generate 4 images for each test prompt and set the same random seed for all methods.
@@
Object Textual Inversion [1gal2022image] TTF SD 1.5 0.199 0.749
Dreambooth [2ruiz2023dreambooth] TTF SD 1.5 0.286 0.772
P+ [23voynov2023p+] TTF SD 1.4 0.244 0.643
Custom Diffusion [6kumari2023multi] TTF SD 1.4 0.307 0.722
NeTI [61alaluf2023neural] TTF SD 1.4 0.283 0.801
SVDiff [8han2023svdiff] TTF SD 1.5 0.282 0.776
Perfusion [7tewel2023key] TTF SD 1.5 0.273 0.691
ELITE [19wei2023elite] PTA SD 1.4 0.292 0.765
BLIP-Diffusion [10li2024blip] PTA SD 1.5 0.292 0.772
IP-Adapter [40ye2023ip] PTA SD 1.5 0.2722 0.825
SSR Encoder [221zhang2024ssr] PTA SD 1.5 0.288 0.792
MoMA [song2024moma] PTA SD 1.5 0.322 0.748
Diptych Prompting [209shin2024large] PTA FLUX 1.0 dev 0.327 0.722
$\lambda$-eclipse [222patel2024lambda] PTA Kandinsky 2.2 0.272 0.824
MS-Diffusion [220personalizationms] PTA SDXL 0.298 0.777
Face CrossInitialization [66pang2023cross] TTF SD 2.1 0.261 0.469
Face2Diffusion [170shiohara2024face2diffusion] PTA SD 1.4 0.265 0.588
SSR Encoder [221zhang2024ssr] PTA SD 1.5 0.233 0.490
FastComposer [54xiao2023fastcomposer] PTA SD 1.5 0.230 0.516
IP-Adapter [40ye2023ip] PTA SD 1.5 0.292 0.462
IP-Adapter [40ye2023ip] PTA SDXL 0.292 0.642
PhotoMaker [28li2023photomaker] PTA SDXL 0.311 0.547
InstantID [37wang2024instantid] PTA SDXL 0.278 0.707
@@ 8.1 Overfitting Problem
To investigate the rationale behind, Compositional Inversion [9zhang2023compositional] observes that the learned token embedding is located in an out-of-distribution area compared to the center distribution formed by pre-trained words. This is also found in another work [66pang2023cross] that the learnable token embeddings deviate significantly from the distribution of the initial embedding. In addition, there is evidence [9zhang2023compositional, 24jia2023taming, 20hao2023vico] suggesting that the unique modifier dominates in the cross-attention layers compared to the other context tokens, leading to the absence of other semantic appearances.
To address this issue, many solutions have been proposed. Most methods discussed in Section 4 contribute to the alleviation of the overfitting problem, such as the exclusion of redundant background, attention manipulation, regularization of the learnable parameters, and data augmentation. However, it has not been solved yet, especially in the cases where the SoI has a non-rigid appearance [9zhang2023compositional] or the context prompt has a similar semantic correlation with the irrelevant elements in the reference [151zhang2024generative]. It is clear that addressing overfitting in PCS is not merely a technical challenge but a necessity for ensuring the practical deployment and scalability of these systems in varied and dynamic real-world environments. Therefore, an effective strategy and robust evaluation metrics are urgently needed to achieve broader adoption and greater satisfaction in practical applications.
@@ 8.2 Trade-off on Subject Fidelity and Text Alignment
The ultimate goal of personalized content synthesis is to create systems that not only render the SoI with high fidelity but also effectively respond to textual prompts. However, achieving excellence in both areas simultaneously presents a notable conflict. Specifically, high subject fidelity typically involves capturing and reproducing detailed and specific features of the SoI. This often requires the model to minimize the reconstruction loss to replicate delicate characteristics accurately. Conversely, text alignment necessitates that the system flexibly adapts the SoI according to varying textual descriptions. These descriptions may suggest changes in pose, expression, environment, or stylistic alterations that do not aim to reconstruct the exact visualization in the reference. As a result, it becomes challenging to achieve flexible adaptation in different contexts while simultaneously pushing the model to capture fine-grained details. To address this inherent conflict, Perfusion [7tewel2023key] proposes to regularize the attention projections by these two items. [35cao2023concept] decouples the conditional guidance into two separate processes, which allows for the distinct handling of subject fidelity and textual alignment. Despite these efforts, there still remains room for further exploration and refinement of this issue. Enhanced model architectures, innovative training methodologies, and more dynamic data handling strategies could potentially provide new pathways to better balance the demands of subject and text fidelity in PCS systems.
@@ 8.4 Multimodal Autoregressive Frameworks
Recent advancements in multimodal autoregressive models present novel solutions for PCS by unifying cross-modal understanding and generation. Models like Emu3 [wang2024emu3] demonstrate that autoregressive architectures can natively handle image-text-video sequences through discrete tokenization and joint transformer training. This paradigm enables seamless integration of user-provided multimodal references (e.g., text descriptions paired with SoI images) while maintaining contextual coherence across generation steps. Besides, this framework natively support for subject editing via multi-round chat, effectively addressing the overfitting limitations commonly observed in diffusion-based models.
@@ 8.5 Interactive Personalization Workflow
The evolution of interactive generation systems has unlocked new frontiers for PCS, particularly through the integration of multi-round interactive generation. This capability allows users to iteratively refine and accurately define the SoI, addressing the challenge of translating vague or complex requirements into precise content generation. For instance, conversational PCS systems like Gemini-2.0-flash [team2023gemini] exemplify this progress, leveraging natural language dialogue to iteratively optimize both subject fidelity and prompt alignment. By enabling users to provide real-time feedback and adjust parameters through chat-like interactions, these systems bridge the gap between abstract intent and concrete outputs, aligning with PCS’s core objective of balancing faithful subject representation with flexible editability.
@@ 9 Conclusion
Textual Inversion [1gal2022image] Object TTF LDM Token Embedding Fine-tuning
@@
DreamBooth [2ruiz2023dreambooth] Object TTF Imagen Diffusion Model Fine-tuning; Regularization Dataset
DreamArtist [68dong2022dreamartist] Object TTF LDM Negative Prompt Fine-tuning
DVAR [69voronov2024loss] Object TTF SD 1.5 Randomness Erasing
HiPer [33han2023highly] Object TTF SD Token Embedding Enhancement
P+ [23voynov2023p+] Object TTF SD 1.4 Token Embedding Enhancement
@@
Unet-finetune [72xiang2023closer] Object TTF SD Parameter-efficient Fine-tuning
Jia et al. [24jia2023taming] Object TTF Imagen Mask-assisted Generation; Regularization
COTI [3yang2023controllable] Object TTF SD 2.0 Data Augmentation
Gradient-Free TI [73fei2023gradient] Object TTF SD Evolutionary Strategy
PerSAM [21zhang2023personalize] Object TTF SD Mask-assisted Generation
DisenBooth [18chen2023disenbooth] Object TTF SD 2.1 Regularization
NeTI [61alaluf2023neural] Object TTF SD 1.4 Token Embedding Enhancement
Prospect [30zhang2023prospect] Object TTF SD 1.4 Token Embedding Enhancement
@@
Break-a-Scene [74avrahami2023break] Object TTF SD 2.1 Attention-based Operation; Attention-based Operation; Mask-assisted Generation
COMCAT [75xiao2023comcat] Object TTF SD 1.4 Attention-based Operation
ViCo [20hao2023vico] Object TTF SD Attention-based Operation; Regularization
OFT [59qiu2024controlling] Object TTF SD 1.5 Finetuning Strategy
LyCORIS [76yeh2023navigating] Object TTF SD 1.5 Attention-based Operation
He et al. [57he2023data] Object TTF SD, SDXL Data Augmentation
DIFFNAT [78roy2023diffnat] Object TTF SD Regularization
@@
MATTE [79agarwal2023image] Object; Style; High-level Semantic TTF SD Regularization
LEGO [45motamed2023lego] Object TTF SD Regularization
Catversion [80zhao2023catversion] Object TTF SD 1.5 Embedding Concatenation
@@
CLiC [81safaee2023clic] Object TTF SD 1.4 Attention-based Operation; Mask-assisted Generation
@@
HiFi Tuner [48wang2023hifi] Object TTF SD 1.4 Attention-based Operation; Mask-assisted Generation; Regularization
InstructBooth [56chae2023instructbooth] Object TTF SD 1.5 Reinforcement Learning
DETEX [85cai2023decoupled] Object TTF SD 1.4 Mask-assisted Generation
DreamDistribution [114zhao2023dreamdistribution] Object TTF SD 2.1 Token Embedding Enhancement
DreamTuner [86hua2023dreamtuner] Object TTF SD Attention-based Operation
Xie et al. [51lu2024object] Object TTF SD Regularization; Mask-assisted Generation
Infusion [157zeng2024infusion] Object TTF SD 1.5 Attention-based Operation
@@
Pair Customization [160jones2024customizing] Object TTF SDXL Disentanglement Approach
DisenDreamer [171chen2024disendreamer] Object TTF SD 2.1 Disentanglement Approach
DreamBooth++++ [219fan2024dreambooth] Object TTF SD Attention-based Operation
CoRe [218wu2024core] Object TTF SD 2.1 Regularization
DCI [216jin2025customized] Object TTF SD Token Embedding Optimization
Re-Imagen [67chen2022re] Object PTA Imagen Retrieval-augmented Paradigm
Versatile Diffusion [109xu2023versatile] Object PTA SD 1.4 Architecture Design
Tuning-Encoder [12gal2023encoder] Object PTA SD Regularization
@@
ELITE [19wei2023elite] Object PTA SD 1.4 Attention-based Operation; Mask-assisted Generation
UMM-Diffusion [71ma2023unified] Object PTA SD 1.5 Reference Feature Injection
SuTI [11chen2024subject] Object PTA Imagen Data Augmentation
InstantBooth [22shi2023instantbooth] Object PTA SD 1.4 Patch Feature Extraction
@@
BLIP-Diffusion [10li2024blip] Object PTA SD 1.5 Data Augmentation; Mask-assisted Generation
@@
Domain-Agnostic [17arar2023domain] Object PTA SD Regularization; Attention-based Operation
IP-Adapter [40ye2023ip] Object PTA SD 1.5 Reference Feature Injection
Kosmos-G [77pan2023kosmos] Object; Face PTA SD 1.5 Multimodal Language Modeling
Kim et al. [47kim2023personalized] Object PTA SD 1.4 Reference Feature Injection
CAFÉ [84zhou2023customization] Object; Face PTA SD 2.0 Multimodal Language Modeling
SAG [99pan2023towards] Object; Face; Style PTA SD 1.5 Gradient-based Guidance
BootPIG [41purushwalkam2024bootpig] Object PTA SD Data Augmentation
CustomContrast [217chen2024customcontrast] Object PTA SD 1.5 Contrastive Learning
MoMA [song2024moma] Object PTA SD 1.5 Multimodal LLM Adapter
@@
Diptych Prompting [209shin2024large] Object; Style PTA FLUX Attention-based Operation; Mask-assisted Generation
@@
StyleDrop [5sohn2024styledrop] Style; Multiple Subjects TTF Muse Data Augmentation
StyleBoost [53park2023styleboost] Style TTF SD 1.5 Regularization Dataset
StyleAligned [98hertz2023style] Style TTF SDXL Regularization
GAL [151zhang2024generative] Style; Object TTF SD 1.5 Data Augmentation
@@
StyleForge [153park2024text] Style TTF SD 1.5 Parameter-efficient Fine-tuning
FineStyle [213zhangfinestyle] Style TTF Muse Mask-assisted Generation
Style-Friendly [211choi2024style] Style TTF FLUX-dev; SD 3.5 SNR Sampler
@@
StyleAdapter [26wang2023styleadapter] Style PTA SD Data Augmentation; Reference Feature Fxtraction
ArtAdapter [42chen2023artadapter] Style PTA SD 1.5 Data Augmentation
ProFusion [55zhou2023enhancing] Face TTF SD 2.0 Fusion Sampling
Celeb Basis [60yuan2023inserting] Face TTF SD Face Feature Representation
Banerjee et al. [39banerjee2023identity] Face TTF SD 1.4 Regularization
@@
Magicapture [90hyung2023magicapture] Face; Multiple Subjects TTF SD 1.5 Attention-based Operation; Mask-assisted Generation; Identity Loss
Concept-centric [35cao2023concept] Face TTF SD 1.5 Modified Classifier-free Guidance
Cross Initialization [66pang2023cross] Face TTF SD 2.1 Regularization
@@
OMG [115kong2024omg] Face; Multiple Subjects TTF SDXL Attention-based Operation; Mask-assisted Generation; Identity Loss
@@
InstantFamily [161kim2024instantfamily] Face; Multiple Subjects; Extra Conditions TTF SD 1.5 Mask-assisted Generation
PersonaMagic [203li2024personamagic] Face TTF SD 1.4 Attention-based Operation
HyperDreamBooth [58ruiz2023hyperdreambooth] Face TTF; PTA SD 1.5 Pretraining and Fast Fine-tuning
Su et al. [49su2023identity] Face PTA LDM; StyleGAN Identity Loss; Multi-task Learning
@@
FastComposer [54xiao2023fastcomposer] Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
Face0 [88valevski2023face0] Face PTA SD 1.4 Fusion Sampling
DreamIdentity [89chen2023dreamidentity] Face PTA SD 2.1 Attention-based Operation
@@
Face-Diffuser [52wang2023high] Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
W-Plus-Adapter [91li2023stylegan] Face PTA SD 1.5; StyleGAN Face Feature Representation
Portrait Diffusion [32liu2023portrait] Face PTA SD 1.5 Mask-assisted Generation
RetriNet [92tang2023retrieving] Face PTA SD Mask-assisted Generation
FaceStudio [31yan2023facestudio] Face PTA SD Attention-based Operation
PVA [27xu2024personalized] Face PTA LDM Mask-assisted Generation
DemoCaricature [94chen2023democaricature] Face; Extra Conditions PTA SD 1.5 Regularization
@@
PhotoMaker [28li2023photomaker] Face PTA SDXL Data Collection; Reference Feature Extraction
Stellar [95achlioptas2023stellar] Face PTA SDXL Evaluation Prompts and Metrics
@@
PortraitBooth [96peng2023portraitbooth] Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
@@
InstantID [37wang2024instantid] Face; Extra Conditions PTA SDXL Multi-feature Injection
ID-Aligner [162chen2024id] Face PTA SD 1.5; SDXL Feedback Learning
@@
MoA [164ostashev2024moa] Face; Multiple Subjects PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
IDAdapter [166cui2024idadapter] Face PTA SD 2.1 Mixed Facial Features
Infinite-ID [169wu2024infinite] Face PTA SDXL Reference Feature Injection
Face2Diffusion [170shiohara2024face2diffusion] Face PTA SD Multi-feature Injection
FreeCure [210cai2024foundation] Face PTA SD 1.5; SDXL Mask-assisted Generation
Omni-ID [205qian2024omni] Face PTA FLUX Face representation enhancement
@@
Custom Diffusion [6kumari2023multi] Multiple Subjects TTF SD 1.4 Parameter-efficient Fine-tuning; Constrained Optimization
Cones [4liu2023cones] Multiple Subjects TTF SD 1.4 Concept Neurons Activation
@@
SVDiff [8han2023svdiff] Multiple Subjects TTF SD Data Augmentation; Attention-based Operation
Perfusion [7tewel2023key] Multiple Subjects TTF SD 1.5 Regularization
@@
Mix-of-Show [46gu2024mix] Multiple Subjects TTF SD 1.5 Attention-based Operation; Mask-assisted Generation
@@
Cones-2 [62liu2023cones] Multiple Subjects TTF SD 2.1 Attention-based Operation; Regularization; Mask-assisted Generation
PACGen [34li2023generate] Multiple Subjects TTF SD 1.4 Mask-assisted Generation
@@
Compositional Inversion [9zhang2023compositional] Multiple Subjects TTF SD Attention-based Operation; Regularization
@@
EM-style [44rahman2024visual] Multiple Subjects TTF SD Attention-based Operation; Mask-assisted Generation
MC2 [154jiang2024mc] Multiple Subjects TTF SD 1.5 Attention-based Operation
MultiBooth [156zhu2024multibooth] Multiple Subjects TTF SD 1.5 Mask-assisted Generation
Matsuda et al. [165matsuda2024multi] Multiple Subjects TTF LDM Mask-assisted Generation
@@
MagicTailor [zhou2024magictailor] Multiple Subjects TTF SD 2.1 Mask-assisted Generation; Attention-based Operation
AnyDoor [14chen2023anydoor] Multiple Subjects PTA SD 2.1 Mask-assisted Generation
@@
Subject-Diffusion [16ma2023subject] Multiple Subjects PTA SD 2.0 Attention-based Operation; Data Augmentation; Mask-assisted Generation
@@
CustomNet [29yuan2023customnet] Multiple Subjects; Extra Conditions PTA SD Multi-condition Integration
@@
MIGC [174zhou2024migc] Multiple Subjects PTA SD Attention-based Operation; Mask-assisted Generation
Emu2 [223sun2024generative] Multiple Subjects PTA SDXL In-Context Learning
SSR-Encoder [221zhang2024ssr] Multiple Subjects PTA SD 1.5 Attention-based Operation